OpenStudy (anonymous):

Prove by contradiction that if a and b are integers and b is odd, then -1 is not a root of f(x)=ax^2+bx+c

OpenStudy (anonymous):

I mean ax^2+bx+a

OpenStudy (anonymous):

0 = a - b + c b = a + c nothing about a or c ?

myininaya (myininaya):

Ok b is odd=>b=2k+1, k is an element of an integer ok so we have \[f(x)=ax^2+(2k+1)x+a\] so its's possible for (-1)=0 well lets see.. \[f(-1)=a(-1)^2+(2k+1)(-1)+a=a-(2k+1)+a\] \[=2a-(2k+1)\] is it possible for 2a-(2k+1)=0 well if a =(2k+1)/2 but this means a is not an integer thus we have received a contradiction

myininaya (myininaya):

no odd number is divisible by 2

myininaya (myininaya):

so its's possible for f(-1)=0*

OpenStudy (anonymous):

why c = a ?

myininaya (myininaya):

thats what he said

OpenStudy (anonymous):

Thank you SO much. Are you a math type of major?

OpenStudy (anonymous):

lol i havent looked on the second thing he wrote :\

OpenStudy (anonymous):

0 = a - b + a b = 2a but 2a is even so b must be even as well